length of portion of the line x=y+2 intersepted by the circle x2+y2=100 is
 

Grade:11

2 Answers

Vikas TU
7 years ago
Just solve the x=y+2 with circle eqn by sy=ubstituting the x value.
and find the x and y coordinates.
then use distnace formulae for calculating the distance..
 
Ajay
7 years ago
The circle has center at O(0, 0) and radius as 10
Ley the the intercept which line cuts the circle be AB and M be mid point of AB
then clearly  OM = distance of line x=y+2 from O(0,0) = sqrt(2)
                 OB = radius = 10
Since Traingle OMB is right angled   MB2 = OB- OM2 =  100-2 = 98
Required length = AB= 2MB = 2sqrt(98)